### v1

# Constraints on SMEFT operators from $Z\to\mu\mu bb$ decay

Abstract
The Standard Model Effective Field Theory (SMEFT) provides a systematic framework to probe indirect effects of heavy new physics via precision measurements. While SMEFT constraints have been extensively studied using purely leptonic $Z$ decays and inclusive $Z$ production, mixed leptonic–hadronic modes remain largely unexplored. In this work, we analyze $Z\to\mu\mu bb$ decays within the SMEFT framework, deriving constraints on dimension-six operators that affect four-fermion interactions between leptons and bottom quarks, as well as ZZ–fermion couplings. Signal and background events are simulated with state-of-the-art Monte Carlo tools, including detector effects such as $b$-tagging, and limits on the relevant Wilson coefficients are extracted using kinematic distributions and a profile likelihood approach. Our results provide complementary constraints to existing SMEFT studies and yield the first process-specific limits on flavor-resolved four-fermion operators involving muons and bottom quarks from $Z$ decays.
preprint: APS/123-QED
I Introduction
The $Z$ boson has long served as a cornerstone of precision tests of the electroweak sector of the Standard Model (SM). At both the LEP and the LHC, measurements of $Z$-boson production and decay have reached a precision that allows indirect probes of physics beyond the SM [1, 2, 3]. In particular, rare and multi-body $Z$-boson decays probe kinematic regimes and operator structures not fully constrained by inclusive observables.
The Standard Model Effective Field Theory (SMEFT) offers a model-independent framework to parameterize the effects of heavy new physics through higher-dimensional operators constructed from SM fields [4, 5, 6]. In recent years, SMEFT interpretations of $Z$-boson observables have attracted considerable attention. Purely leptonic decay modes, such as $Z\to 4\ell$, provide stringent constraints on electroweak and four-lepton operators, while complementary information has been obtained from SMEFT analyses of inclusive $Z$ production and ZZ+jets, probing quark electroweak couplings and dipole interactions [7, 8, 9, 10].
Despite this progress, $Z$-boson decays into mixed leptonic–hadronic final states remain relatively unexplored in the SMEFT context. Such channels are sensitive to a distinct set of operators, in particular four-fermion contact interactions involving both leptons and quarks, as well as flavor-dependent modifications of ZZ–fermion couplings [11]. Among these, the decay $Z\to\mu\mu bb$ is of special interest. The presence of bottom quarks enhances sensitivity to operators coupling to the third quark generation, while the dimuon final state allows for precise reconstruction of the leptonic kinematics. At the same time, the analysis poses experimental challenges due to sizable QCD backgrounds and the reliance on heavy-flavor tagging [12, 13].
Figure 1: Representative Feynman diagram contributing to the process $Z\to\mu\mu bb$ in the Standard Model, where the four-body final state arises from final-state radiation of an off-shell $Z^{\ast}/\gamma^{\ast}$.
The diagram shown in FIG. 1 illustrates the characteristic topology of the $Z\to\mu\mu bb$ decay in the Standard Model, where the four-body final state arises predominantly from final-state radiation of an off-shell electroweak gauge boson. Compared to two-body $Z$ decays, this topology probes a richer kinematic structure and allows access to contact interactions that are absent at leading order in simpler final states. In the SMEFT framework, this process is particularly sensitive to four-fermion operators involving muons and bottom quarks, which generate direct $\mu\mu bb$ interactions at tree level. Such contributions interfere with the SM amplitudes mediated by off-shell $Z^{\ast}/\gamma^{\ast}$ exchange, leading to characteristic deviations in both the total rate and differential distributions. As a result, the $Z\to\mu\mu bb$ channel provides a complementary, process-specific probe of lepton–quark operator structures that are only weakly constrained by inclusive $Z$ observables or purely leptonic decay modes.
In this work, we perform a dedicated SMEFT analysis of the decay $Z\to\mu\mu bb$ and derive constraints on dimension-six operators contributing to this process. We focus on operators that induce four-fermion interactions of the $\ell\ell bb$ type and on operators that modify the $Z$ couplings to leptons and bottom quarks. Signal and background processes are simulated using state-of-the-art Monte Carlo tools, and detector-level effects are included [14, 15, 16]. Limits on the WCs are obtained from differential distributions using a profile likelihood approach [17].
This study provides a complementary probe of new physics effects in the $Z$-boson sector, extending existing SMEFT constraints beyond purely leptonic final states. It demonstrates the potential of mixed leptonic–hadronic $Z$ decays to access operator directions that are weakly constrained by current analyses and motivates their inclusion in future global SMEFT fits [18].
II Effective field theory description
In the absence of direct evidence for new resonances, the effects of physics beyond the Standard Model (SM) can be systematically described using an effective field theory approach [4, 5, 6]. Its low-energy effects can be encoded in higher-dimensional operators constructed from SM fields and respecting the SM gauge symmetries.
Within the SMEFT framework, the Lagrangian is organized as an expansion in inverse powers of the new physics scale $\Lambda$,
$\mathcal{L}_{\mathrm{SMEFT}}=\mathcal{L}_{\mathrm{SM}}+\sum_{i}\frac{C_{i}}{\Lambda^{2}}\mathcal{O}_{i}+\mathcal{O}\left(\frac{1}{\Lambda^{4}}\right),$ (1)
where $\mathcal{O}_{i}$ are dimension-six operators and $C_{i}$ the corresponding WCs [10, 11]. In this work, we truncate the expansion at dimension six contributions [19, 20].
The decay $Z\to\mu\mu bb$ receives contributions from several classes of dimension-six operators. First, four-fermion operators involving leptons and bottom quarks generate direct contact interactions contributing to the four-body final state at tree level [7, 8]. Schematically, such operators can be written as $(\bar{\ell}\Gamma\ell)(\bar{q}\Gamma q)$ and induce amplitudes that interfere with the SM contributions mediated by off-shell electroweak bosons.
Second, operators that modify the couplings of the $Z$ boson to fermions lead to indirect contributions to the decay process. After electroweak symmetry breaking, these operators generate shifts in the effective ZZ–lepton and ZZ–quark couplings, which affect both the production and decay of the $Z$ boson [21, 18]. These effects propagate into the $Z\to\mu\mu bb$ final state through the modified electroweak vertices.
In the numerical analysis, we select a set of six dimension-six operators that contribute to this decay channel and study their impact individually. The operators are implemented using the SMEFTsim_general_MwScheme_UFO model [6, 22], which is compatible with MadGraph5_aMC@NLO [14] and allows event generation and reweighting at leading order in a consistent electroweak input scheme.
In this analysis, we retain both the interference and quadratic terms in the WCs when evaluating SMEFT effects [7]. The dependence of physical observables on a given WC is therefore parameterized by a quadratic function, which can be efficiently determined using the reweighting technique [14]. This approach allows us to explore the parameter space of the WCs without generating independent event samples for each point.
The validity of the SMEFT expansion relies on the assumption that the typical energy scale of the process remains below the cutoff scale $\Lambda$ [19, 10]. In practice, the absence of events with extremely large $m_{\mu\mu bb}$ or jet transverse momenta indicates that the analysis is dominated by energy scales safely below the assumed new physics cutoff.
III Event generation and simulation
All processes are generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8 [15]. Detector-level effects are simulated using Delphes [16] with an CMS-like detector configuration. Unless stated otherwise, proton–proton collisions at a center-of-mass energy of $\sqrt{s}=13~\mathrm{TeV}$ are assumed throughout this analysis.
The signal process considered in this work is the four-body decay of the Z boson into a dimuon pair and a bottom–antibottom quark pair: $Z\to\mu\mu bb$. At leading order, the four-body final state arises predominantly from final-state radiation or off-shell electroweak boson exchange [7]. Representative Feynman diagrams contributing to the signal process are shown in Fig. 1.
Signal events are generated at parton level using MadGraph5_aMC@NLO. Basic generator-level cuts are applied to ensure numerical stability and reflect the kinematic acceptance of the detector: $p_{T}^{\ell}>10~\mathrm{GeV}$, $|\eta^{\ell}|<2.5$ for charged leptons, and $p_{T}^{q}>10~\mathrm{GeV}$, $|\eta^{q}|<2.5$ for quarks. The resulting leading-order cross section for the signal process is $\sigma(Z\to\mu^{+}\mu^{-}b\bar{b})=0.750~\mathrm{fb}$, where the quoted uncertainty corresponds to the Monte Carlo integration error.
The reconstructed kinematic distributions shown in Fig. 2 provide an important validation of the signal modeling after detector effects are taken into account. In Fig. 2(a), the invariant mass distribution of the $b\bar{b}$ system is considerably broader than that of dimuon system, as expected from jet energy resolution effects and QCD radiation. Nevertheless, the combined four-body invariant mass $m_{\mu\mu bb}$ displays a clear enhancement around $m_{Z}$, indicating that the signal topology can be reliably reconstructed at detector level despite the presence of hadronic final states.
Fig. 2(b) shows the transverse momentum spectra of the reconstructed muons and jets. The muon $p_{T}$ distribution is relatively soft, characteristic of a $Z$-boson decay, while the jet $p_{T}$ spectrum features a longer tail driven by QCD dynamics. The typical energy scales involved remain well below the assumed SMEFT cutoff, supporting the consistency of the SMEFT interpretation within the selected phase space.
((a))
((b))
Figure 2: Detector-level kinematic distributions for the process $Z\to\mu\mu bb$ obtained from Delphes simulation.
The dominant background arises from top-quark pair production $pp\to t\bar{t}$, with both top quarks decaying semileptonically, $t\to b\mu^{+}\nu_{\mu}$ and $\bar{t}\to\bar{b}\mu^{-}\bar{\nu}_{\mu}$ [14, 15]. Due to its large production cross section, this process remains a significant background even after event selection.
Additional backgrounds include diboson production, such as $pp\to ZZ$, $Z\to\mu\mu$, $Z\to bb$ as well as associated production of a Z boson with a Higgs boson: $pp\to ZH$, $Z\to\mu^{+}\mu^{-}$, $H\to b\bar{b}$. Although their cross sections are smaller than that of $t\bar{t}$ production, these processes closely resemble the signal topology and are therefore included explicitly [16, 14].
The Drell–Yan production of a $Z$ boson in association with jets: $pp\to Z+\text{jets}$, constitutes another important background due to its very large cross section. Events with light-flavor jets can enter the signal region through jet misidentification, while events with heavy-flavor jets provide an irreducible background. To model this background accurately, samples with up to two additional jets at matrix-element level are generated and merged with the parton shower using the MLM matching scheme [23, 24].
Subleading backgrounds from electroweak processes involving photons are also considered, including $Z\gamma$, $\gamma\gamma$, and mixed $\gamma Z$ production, where the photon converts into a fermion pair.
A summary of all signal and background processes and their corresponding leading-order cross sections is given in TABLE 1.
Process Cross section [pb]
$pp\to Z,\ Z\to\mu^{+}\mu^{-}b\bar{b}$ $7.50\times 10^{-4}$
$pp\to t\bar{t}$, $t\to b\mu^{+}\nu_{\mu},\bar{t}\to\bar{b}\mu^{-}\bar{\nu}_{\mu}$ $4.38$
$pp\to ZZ$, $Z\to\mu^{+}\mu^{-}$, $Z\to b\bar{b}$ $5.54\times 10^{-2}$
$pp\to Z+\text{jets}$ $(0,1,2)$ $9.33\times 10^{2}$
$pp\to ZH$, $Z\to\mu^{+}\mu^{-}$, $H\to b\bar{b}$ $1.03\times 10^{-2}$
$pp\to Z\gamma$, $\gamma\to b\bar{b}$ $5.54\times 10^{-3}$
$pp\to\gamma\gamma$, $\gamma\to\mu^{+}\mu^{-},\ \gamma\to b\bar{b}$ $6.70\times 10^{-5}$
Table 1: Leading-order cross sections for the signal and background processes considered in this analysis. The quoted uncertainties correspond to Monte Carlo integration errors. The merged ZZ+jets sample, combining matrix-element calculations with up to two additional jets, is used in the final analysis, while the individual samples are shown for reference.
All generated parton-level events are passed to Pythia8 for parton showering and hadronization. For the $Z+\text{jets}$ background, matrix-element calculations with up to two additional jets are merged with the parton shower using the MLM matching scheme [23, 24]. This procedure avoids double counting between matrix-element and parton-shower emissions and provides a more reliable description of multi-jet final states.
Detector effects are simulated with Delphes, using a CMS-like detector configuration. Jets are identified as $b$-jets using the built-in $b$-tagging algorithm, and a fixed working point corresponding to a discriminator threshold of $\text{BTag}>0.5$ is applied. This working point yields an average $b$-tagging efficiency of approximately $70\%$ for true $b$-jets, with mis-tag rates of about $10\%$ for charm jets and $1\%$ for light-flavor jets, consistent with typical CMS performance [25]. The requirement of two $b$-tagged jets significantly suppresses backgrounds from light-flavor and charm jet production, in particular from Z+Z+jets events.
IV Event Reconstruction and Selection Strategy
Events are reconstructed using the detector-level objects provided by Delphes. Muon candidates are reconstructed within the pseudorapidity range $|\eta|<2.5$, corresponding to the coverage of the central tracking system and muon chambers in LHC experiments [26, 27]. Jets are reconstructed using the anti-$k_{T}$ algorithm with a radius parameter $R=0.4$ [28].
Bottom-quark jets are identified using the $b$-tagging information stored in the Delphes output [16, 29]. A jet is considered a $b$-jet if its $b$-tag discriminator exceeds a threshold of $0.5$, corresponding to a typical medium working point used in CMS analyses. Jets are identified as $b$-jets using the $b$-tagging discriminator provided by Delphes, adopting a medium working point.
To emulate the trigger requirements of LHC experiments, events are required to satisfy at least one of the muon-based trigger conditions listed in TABLE 2. These triggers are motivated by standard CMS single- and double-muon triggers used in Run II analyses [30]. The single-muon trigger ensures high efficiency for events containing a high-$p_{T}$ muon, while the di-muon trigger retains sensitivity to events in which both muons have moderate transverse momenta. Only isolated muons are considered for trigger selection to suppress backgrounds from heavy-flavor decays and hadronic activity.
Trigger requirement Threshold [GeV]
Single-muon trigger $p_{T}(\mu)>25$
Double-muon trigger (leading) $p_{T}(\mu_{1})>18$
Double-muon trigger (subleading) $p_{T}(\mu_{2})>7$
Table 2: Muon trigger requirements used in the analysis. Only isolated muons are considered for trigger selection.
Events passing the trigger requirements are subjected to a set of preselection criteria designed to suppress reducible backgrounds while maintaining high signal efficiency. The preselection cuts are summarized in TABLE 3.
At least two muon candidates are required per event. The subleading muon is required to satisfy $p_{T}>10~\mathrm{GeV}$ to ensure reliable reconstruction and identification efficiency. Muon isolation is imposed by requiring a minimum separation $\Delta R(\mu,\text{jet})>0.4$ between each selected muon and any reconstructed jet [28]. This reduces backgrounds from semileptonic heavy-flavor decays.
To suppress backgrounds containing genuine missing transverse energy, such as top-quark pair production, events are required to have $\mathrm{MET}<30~\mathrm{GeV}$. This exploits the absence of neutrinos in the signal process $Z\to\mu\mu bb$.
Events are required to contain at least two reconstructed jets with $p_{T}>25~\mathrm{GeV}$, among which at least two must be identified as $b$-jets. This reflects the expected presence of a $b\bar{b}$ pair in the signal final state and significantly suppresses backgrounds from $Z+$ light-jets production.
Angular separation requirements are imposed to ensure well-resolved final-state objects. In particular, the separation between the two selected $b$-jets must satisfy $\Delta R(b,b)>0.4$, and the separation between the two leading muons must satisfy $\Delta R(\mu,\mu)>0.4$. Additionally, all muon–$b$-jet pairs must satisfy $\Delta R(\mu,b)>0.4$ [28]. These requirements reduce contamination from collinear final-state radiation and overlapping reconstruction effects.
Selection Requirement
Number of muons $\geq 2$
Muon $p_{T}$ (subleading) $10~\mathrm{GeV}$
Muon isolation $\Delta R(\mu,\text{jet})>0.4$
Missing transverse energy $\mathrm{MET}<30~\mathrm{GeV}$
Number of jets $\geq 2$
Jet $p_{T}$ $25~\mathrm{GeV}$
$b$-tagging $\geq 2$ $b$-jets
$\Delta R(b,b)$ >0.4>0.4
$\Delta R(\mu,\mu)$ >0.4>0.4
$\Delta R(\mu,b)$ >0.4>0.4
Table 3: Summary of event selection criteria applied in the analysis.
The selection criteria are designed to preserve the characteristic kinematics of the $Z\to\mu\mu bb$ decay while suppressing backgrounds with genuine missing energy or misidentified heavy-flavor jets. For events passing all selection requirements, the invariant mass of the $\mu\mu bb$ system is reconstructed using the four-momenta of the two leading isolated muons and the two selected $b$-jets. The reconstructed invariant mass $m_{\mu\mu bb}$, serves as the primary observable for signal extraction and the subsequent SMEFT analysis [7].
The signal region is defined as $80~\mathrm{GeV}<m_{\mu\mu bb}<115~\mathrm{GeV}$, centered around the $Z$ boson mass. This window retains the majority of the signal while efficiently suppressing non-resonant backgrounds. This window is chosen to maximize sensitivity to SMEFT-induced rate deviations while minimizing contamination from non-resonant backgrounds.
FIG. 3 compares the reconstructed invariant mass distributions of the $\mu\mu bb$ system for the signal and the dominant background processes after the full event selection. In FIG. 3(a), the distribution is shown over a wide mass range from 70 to 250 GeV, with the signal contribution scaled by a factor of 100 for visibility. Over this range, the spectrum is dominated by the ZZ+jets and $t\bar{t}$ backgrounds, reflecting their large production cross sections.
((a))
((b))
Figure 3: Detector-level invariant mass distributions for $\mu\mu bb$ from various processes, obtained from Delphes simulation. The first panel shows the full mass spectrum with the signal scaled for better visibility, while the second panel zooms into the signal region.
FIG. 3(b) focuses on the signal region defined by $80<m_{\mu\mu bb}<115$ GeV, which is centered around the $Z$ boson mass. In this window, the signal contribution becomes clearly distinguishable, while the $t\bar{t}$ background is strongly suppressed by the missing transverse energy requirement. The remaining backgrounds, including $ZA$ and $ZH$ production, exhibit similar kinematic features but are subleading in rate. This demonstrates that the chosen invariant mass window provides an optimal balance between signal efficiency and background rejection, forming a robust basis for the subsequent SMEFT analysis. The clear localization of the signal contribution within the $Z$-mass window demonstrates that the invariant mass $m_{\mu\mu bb}$ provides a powerful discriminator for isolating SMEFT effects in this channel. This observable is therefore chosen as the primary input for the subsequent reweighting and limit-setting procedure.
V SMEFT Reweighting and Constraining
The impact of dimension-six operators on the process $Z\to\mu\mu bb$ is evaluated within the SMEFT framework using an event reweighting technique [31, 32]. In this approach, the effect of each WC on the cross section and kinematic distributions is obtained by reweighting SM Monte Carlo events. The total event weight includes the Standard Model contribution as well as linear and quadratic corrections from dimension-six operators, allowing an efficient exploration of the WCs parameter space without regenerating new event samples.
SMEFT effects are incorporated using the SMEFTsim_general_MwScheme_UFO model from the SMEFTsim 3.0 framework [6, 22], which provides a complete and gauge-invariant implementation of dimension-six operators in the $M_{W}$ input scheme. The $M_{W}$ scheme is adopted to consistently account for electroweak input-parameter shifts induced by dimension-six operators. For each generated event, a set of precomputed weights corresponding to different values of $\theta$ is stored and propagated through parton showering, hadronization, and detector simulation, ensuring a consistent treatment of SM and SMEFT contributions [15, 16].
After applying the full event selection, the expected number of signal events is expressed as a quadratic function of the WCs parameter $\theta\equiv C_{i}/\Lambda^{2}$,
$N(\theta)=c+b\,\theta+a\,\theta^{2},$ (2)
where $c$ denotes the Standard Model (SM) expectation, $b$ corresponds to the interference between the SM and dimension-six amplitudes, and $a$ encodes the pure quadratic SMEFT contribution [7]. The coefficients $a$, $b$, and $c$ are determined by fitting the event yields obtained from reweighted samples generated at several fixed values of $\theta$.
Constraints on the WCs are derived using a likelihood-based statistical approach [17]. An Asimov dataset is constructed by assuming the observed number of events to be equal to the SM prediction, $N_{\mathrm{obs}}=N_{\mathrm{SM}}\equiv c$. The Poisson likelihood function is then given by
$\mathcal{L}(\theta)=\frac{[N(\theta)]^{N_{\mathrm{obs}}}e^{-N(\theta)}}{N_{\mathrm{obs}}!}.$ (3)
The test statistic is defined as the likelihood ratio
$q(\theta)=-2\ln\frac{\mathcal{L}(\theta)}{\mathcal{L}(0)}=2\left[N(\theta)\ln\frac{N(\theta)}{N_{\mathrm{SM}}}+N_{\mathrm{SM}}-N(\theta)\right],$ (4)
which, under the assumptions of Wilks’ theorem, follows a $\chi^{2}$ distribution with one degree of freedom [33]. The 95% confidence level (C.L.) interval on $\theta$ is obtained by requiring $q(\theta)=3.84$.
The resulting constraint is quoted both in terms of the WCs, $\theta_{95}=C_{i}/\Lambda^{2}$, and in terms of the relative deviation of the expected event yield with respect to the SM prediction,
$R_{95}\equiv\frac{N(\theta_{95})}{N_{\mathrm{SM}}}.$ (5)
This procedure is applied independently to each dimension-six operator, with all other WCs set to zero, providing a measure of the sensitivity of the $Z\to\mu\mu bb$ channel to physics beyond the Standard Model within the SMEFT framework. In this analysis, a flavor-specific setup is adopted, in which the WCs are defined for individual fermion generations. In particular, operators involving second-generation leptons ($\ell_{2}$) and third-generation quarks ($q_{3}$) are considered explicitly. This choice allows a direct probe of flavor-resolved SMEFT effects in the $Z\to\mu\mu bb$ final state, beyond the flavor-universal assumptions commonly employed in global SMEFT fits.
FIG. 4 shows the dependence of the expected event yield in the signal region on the WCs of the six dimension-six operators considered in this analysis. For each operator, the ratio of the SMEFT prediction to the SM expectation is fitted with a quadratic function, reflecting the structure of the squared matrix element, which includes both interference and pure SMEFT contributions.
The observed quadratic behavior confirms the validity of the reweighting procedure and illustrates the relative importance of linear and quadratic terms for different operator classes. Operators modifying the ZZ–fermion couplings exhibit sizable interference with the SM amplitude, leading to asymmetric constraints around the SM point. In contrast, certain four-fermion operators show a more symmetric dependence, indicating a dominant contribution from the quadratic term. This behavior highlights the particular sensitivity of this channel to four-fermion operators involving bottom quarks.
The resulting 95% C.L. intervals on the WCs obtained from this procedure are summarized in Table 4. The constraints obtained in this analysis can be compared with existing global SMEFT interpretations from the CMS Collaboration [34]. In the CMS combined analysis, stringent limits are reported on operators that modify the couplings of electroweak gauge bosons to fermions, such as $C_{H\ell}^{(1,3)}$ and $C_{Hq}^{(1,3)}$, under flavor-universal or top-specific assumptions. These limits, typically at the level of $\mathcal{O}(10^{-3})$$\mathcal{O}(10^{-2})~\mathrm{TeV}^{-2}$ at 95% C.L., are primarily driven by electroweak precision observables, Higgs measurements, and top-quark processes.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 4: Quadratic SMEFT fits for the six operators under study. Each panel shows the fitted yield as a function of the WCs for the corresponding operator, with points representing simulated events and the red line indicating the quadratic fit. The legend shows the operator name following the SMEFTsim 3.0 conventions.
In contrast, the present analysis focuses on flavor-resolved WCs involving second-generation leptons and third-generation quarks. The constraints derived for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ can therefore be directly interpreted as channel-specific probes of the same operator classes considered in the CMS global fit, but obtained from a single decay mode and a cut-based analysis.
The constraints obtained here for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ are found to be of the same order of magnitude as those from the CMS global fit, despite being derived from a single decay channel and a cut-based analysis. For the four-fermion operators $C_{\ell q}^{(1,3),2233}$, which describe contact interactions between second-generation leptons and third-generation quarks, no direct constraints are currently available from global SMEFT fits, where such flavor-specific four-fermion operators are typically neglected or marginalized. The limits presented in this work therefore provide the first process-specific constraints on these operators from $Z$-boson decays. These results demonstrate the unique sensitivity of the $Z\to\mu\mu bb$ final state to four-fermion interactions involving bottom quarks, which are difficult to access in inclusive or flavor-universal analyses. The projected HL-LHC dataset, with an integrated luminosity of 3000 fb-1, further improves the constraints and providing a flavor-resolved probe complementary to existing CMS global analyses. In particular, the HL-LHC projections confirm the unique potential of the $Z\to\mu\mu bb$ channel to constrain four-fermion interactions involving bottom quarks.
$C$ 138 fb-1 3000 fb-1
$C_{\ell q}^{(1),2233}$ [0.023, 0.014][-0.023,\;0.014] [0.005, 0.003][-0.005,\;0.003]
$C_{\ell q}^{(3),2233}$ [0.023, 0.014][-0.023,\;0.014] [0.005, 0.003][-0.005,\;0.003]
$C_{Hq}^{(1),33}$ [0.029, 0.026][-0.029,\;0.026] [0.006, 0.006][-0.006,\;0.006]
$C_{Hq}^{(3),33}$ [0.030, 0.026][-0.030,\;0.026] [0.006, 0.006][-0.006,\;0.006]
$C_{H\ell}^{(1),22}$ [0.025, 0.009][-0.025,\;0.009] [0.005, 0.002][-0.005,\;0.002]
$C_{H\ell}^{(3),22}$ [0.028, 0.025][-0.028,\;0.025] [0.006, 0.005][-0.006,\;0.005]
Table 4: 95% C.L. allowed regions for $C/\Lambda^{2}$ [TeV-2].
VI Conclusions
In this work, we have presented a phenomenological study of the $Z\to\mu\mu bb$ final state as a probe of physics beyond the Standard Model within the framework of the Standard Model Effective Field Theory. The analysis is entirely based on Monte Carlo simulations and aims at assessing the sensitivity of this channel to a selected set of dimension-six SMEFT operators.
Signal and relevant background processes were generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia, and a fast detector simulation with Delphes. A realistic event selection strategy was implemented, including trigger requirements, lepton isolation, $b$-jet identification, and kinematic separation cuts. After event reconstruction, the invariant mass distribution of the $\mu\mu bb$ system, together with transverse momentum spectra of the final-state objects, was used to validate the signal modeling and background suppression.
To quantify possible deviations from the Standard Model, we employed the SMEFT reweighting technique using a UFO implementation of dimension-six operators. For each operator under consideration, event weights corresponding to multiple values of the WC were extracted at the generator level and propagated through the full analysis chain. The total event yield after selection was shown to follow a quadratic dependence on the WC, reflecting the interference and pure SMEFT contributions to the cross section.
A likelihood-based statistical analysis was performed by constructing a test statistic based on the expected event yields. Assuming the Standard Model prediction as the null hypothesis, we derived 95% confidence level constraints on the WCs by identifying the parameter values corresponding to the critical value of the test statistic. This procedure provides a transparent and robust estimate of the sensitivity of the $\mu\mu bb$ channel to SMEFT effects, in the absence of experimental systematic uncertainties associated with real data.
The results demonstrate that the $Z\to\mu\mu bb$ final state constitutes a viable and complementary probe of electroweak and Higgs-related effective operators. Although the present study is limited to leading-order simulations and an inclusive cut-based analysis, it establishes a solid baseline for future improvements. Possible extensions include the incorporation of higher-order corrections, a differential shape analysis, the inclusion of systematic uncertainties, and a global fit combining multiple final states.
Overall, this work highlights the potential of precision studies of multi-body $Z$-boson decay final states within the SMEFT framework and demonstrates that mixed leptonic–hadronic channels can provide unique and complementary information on flavor-resolved new physics interactions.
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grants No. 12325504.

### v3

@@
 Abstract
-The Standard Model Effective Field Theory (SMEFT) provides a systematic framework to probe indirect effects of heavy new physics via precision measurements.
-While SMEFT constraints have been extensively studied using purely leptonic $Z$ decays and inclusive $Z$ production, mixed leptonic–hadronic modes remain largely unexplored.
-In this work, we analyze $Z\to\mu\mu bb$ decays within the SMEFT framework, deriving constraints on dimension-six operators that affect four-fermion interactions between leptons and bottom quarks, as well as ZZ–fermion couplings.
-Signal and background events are simulated with state-of-the-art Monte Carlo tools, including detector effects such as $b$-tagging, and limits on the relevant Wilson coefficients are extracted using kinematic distributions and a profile likelihood approach.
-Our results provide complementary constraints to existing SMEFT studies and yield the first process-specific limits on flavor-resolved four-fermion operators involving muons and bottom quarks from $Z$ decays.
+The Standard Model Effective Field Theory (SMEFT) provides a systematic framework to parametrize indirect effects of heavy new physics in precision measurements.
+In this work, we study the $\mu^{+}\mu^{-}b\bar{b}$ final state in the reconstructed $Z$-pole region and derive constraints on selected dimension-six SMEFT operators involving second-generation leptons and third-generation quarks.
+Signal and background processes are simulated at leading order using standard Monte Carlo tools, followed by parton showering and fast detector simulation with simplified $b$-tagging effects.
+We focus on four-fermion operators involving leptons and bottom quarks, as well as operators modifying effective ZZ–fermion couplings.
+The dependence of the selected event yield on individual Wilson coefficients is obtained using SMEFT reweighting and parametrized including both interference and quadratic dimension-six contributions.
+Expected constraints are derived with an Asimov likelihood-ratio approach for integrated luminosities of $138~\mathrm{fb}^{-1}$ and $3000~\mathrm{fb}^{-1}$.
+The results provide channel-specific constraints on flavor-resolved SMEFT interactions in mixed leptonic–hadronic $Z$-pole final states and offer a useful reference for future analyses including higher-order and systematic effects.
 preprint: APS/123-QED
@@
 In recent years, SMEFT interpretations of $Z$-boson observables have attracted considerable attention.
-Purely leptonic decay modes, such as $Z\to 4\ell$, provide stringent constraints on electroweak and four-lepton operators, while complementary information has been obtained from SMEFT analyses of inclusive $Z$ production and ZZ+jets, probing quark electroweak couplings and dipole interactions [7, 8, 9, 10].
+Purely leptonic decay modes, such as $Z\to 4\ell$, provide stringent constraints on electroweak and four-lepton operators, while complementary information has been obtained from SMEFT analyses of inclusive $Z$ production and ZZ+jets, probing quark electroweak couplings and dipole interactions [7, 8].
 Despite this progress, $Z$-boson decays into mixed leptonic–hadronic final states remain relatively unexplored in the SMEFT context.
 Such channels are sensitive to a distinct set of operators, in particular four-fermion contact interactions involving both leptons and quarks, as well as flavor-dependent modifications of ZZ–fermion couplings [9].
-Among these, the decay $Z\to\mu\mu bb$ is of special interest.
+Among these, the $\mu^{+}\mu^{-}b\bar{b}$ final state in the $Z$-pole region is of special interest.
 The presence of bottom quarks enhances sensitivity to operators coupling to the third quark generation, while the dimuon final state allows for precise reconstruction of the leptonic kinematics.
 At the same time, the analysis poses experimental challenges due to sizable QCD backgrounds and the reliance on heavy-flavor tagging [10, 11].
-Figure 1: Representative Feynman diagram contributing to the process $Z\to\mu\mu bb$ in the Standard Model, where the four-body final state arises from final-state radiation of an off-shell $Z^{\ast}/\gamma^{\ast}$.
+Figure 1: Representative Feynman diagram contributing to the $\mu^{+}\mu^{-}b\bar{b}$ final state near the $Z$ pole in the Standard Model, where the four-body final state may arise from final-state radiation of an off-shell $Z^{\ast}/\gamma^{\ast}$.
 The diagram shown in FIG.
@@
 As a result, the $Z\to\mu\mu bb$ channel provides a complementary, process-specific probe of lepton–quark operator structures that are only weakly constrained by inclusive $Z$ observables or purely leptonic decay modes.
-In this work, we perform a dedicated SMEFT analysis of the decay $Z\to\mu\mu bb$ and derive constraints on dimension-six operators contributing to this process.
-We focus on operators that induce four-fermion interactions of the $\ell\ell bb$ type and on operators that modify the $Z$ couplings to leptons and bottom quarks.
-Signal and background processes are simulated using state-of-the-art Monte Carlo tools, and detector-level effects are included [14, 15, 16].
-Limits on the WCs are obtained from differential distributions using a profile likelihood approach [17].
-This study provides a complementary probe of new physics effects in the $Z$-boson sector, extending existing SMEFT constraints beyond purely leptonic final states.
-It demonstrates the potential of mixed leptonic–hadronic $Z$ decays to access operator directions that are weakly constrained by current analyses and motivates their inclusion in future global SMEFT fits [18].
+The analysis is based on leading-order Monte Carlo simulations combined with parton showering and fast detector simulation.
+Constraints on the selected Wilson coefficients are obtained with an Asimov likelihood-ratio method by varying one operator at a time.
+Within this setup, the study provides channel-specific constraints on flavor-resolved SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state.
+The results also provide a useful starting point for future refinements including higher-order corrections, systematic uncertainties, electroweak interference effects, and multi-operator SMEFT interpretations.
 II Effective field theory description
@@
 $\mathcal{L}_{\mathrm{SMEFT}}=\mathcal{L}_{\mathrm{SM}}+\sum_{i}\frac{C_{i}}{\Lambda^{2}}\mathcal{O}_{i}+\mathcal{O}\left(\frac{1}{\Lambda^{4}}\right),$ (1)
-where $\mathcal{O}_{i}$ are dimension-six operators and $C_{i}$ the corresponding WCs [10, 11]. In this work, we truncate the expansion at dimension six contributions [19, 20].
-The decay $Z\to\mu\mu bb$ receives contributions from several classes of dimension-six operators.
-First, four-fermion operators involving leptons and bottom quarks generate direct contact interactions contributing to the four-body final state at tree level [7, 8].
+where $\mathcal{O}_{i}$ are dimension-six operators and $C_{i}$ the corresponding WCs [8, 9].
+In this work, we focus on the effects of dimension-six operators.
+Both the interference between the SM and dimension-six amplitudes and, in the baseline numerical results, the quadratic contribution from dimension-six amplitudes are considered.
+The latter is formally of order $1/\Lambda^{4}$ and should therefore be interpreted with the usual SMEFT truncation caveat.
+For a generic event yield, the SMEFT dependence can be written as
+$N(\vec{C})=N_{\mathrm{SM}}+\sum_{i}\frac{C_{i}}{\Lambda^{2}}N_{i}^{\mathrm{int}}+\sum_{i,j}\frac{C_{i}C_{j}}{\Lambda^{4}}N_{ij}^{\mathrm{quad}},$ (2)
+where $N_{i}^{\mathrm{int}}$ denotes the interference between the SM amplitude and the amplitude induced by a single dimension-six operator, while $N_{ij}^{\mathrm{quad}}$ denotes the quadratic contribution from dimension-six amplitudes.
+In the numerical analysis below, one operator is varied at a time, so that Eq.
+(2) reduces to a quadratic dependence on a single parameter.
+The resulting limits are therefore individual expected constraints and should not be interpreted as fully marginalized SMEFT constraints.
+The selected $\mu^{+}\mu^{-}b\bar{b}$ final state near the $Z$ pole receives contributions from several classes of dimension-six operators.
+First, four-fermion operators involving leptons and bottom quarks generate direct contact interactions contributing to the four-body final state at tree level [7].
 Schematically, such operators can be written as $(\bar{\ell}\Gamma\ell)(\bar{q}\Gamma q)$ and induce amplitudes that interfere with the SM contributions mediated by off-shell electroweak bosons.
@@
 These effects propagate into the $Z\to\mu\mu bb$ final state through the modified electroweak vertices.
+Throughout the event generation and reweighting procedure, the total $Z$-boson width is kept fixed to its SM value.
+Operators such as $C_{H\ell}^{(1,3)}$ and $C_{Hq}^{(1,3)}$ can in general modify the total and partial $Z$ widths.
+The expected constraints reported in this work should therefore be interpreted under the fixed-width assumption.
+Including SMEFT-induced width variations would require a global treatment of electroweak precision observables and is left for future work.
 In the numerical analysis, we select a set of six dimension-six operators that contribute to this decay channel and study their impact individually.
-The operators are implemented using the SMEFTsim_general_MwScheme_UFO model [6, 22], which is compatible with MadGraph5_aMC@NLO [14] and allows event generation and reweighting at leading order in a consistent electroweak input scheme.
-In this analysis, we retain both the interference and quadratic terms in the WCs when evaluating SMEFT effects [7].
-The dependence of physical observables on a given WC is therefore parameterized by a quadratic function, which can be efficiently determined using the reweighting technique [14].
-This approach allows us to explore the parameter space of the WCs without generating independent event samples for each point.
+The operators are implemented using the SMEFTsim_general_MwScheme_UFO model [14], which is compatible with MadGraph5_aMC@NLO [15] and allows event generation and reweighting at leading order in a consistent electroweak input scheme.
+In the baseline numerical results, both interference and quadratic dimension-six terms are retained in the event yield.
+This convention follows the structure of the squared matrix element used in the reweighting procedure and is particularly relevant for operators with suppressed interference.
+The associated EFT truncation caveat is discussed above and is taken into account when interpreting the quoted constraints.
 The validity of the SMEFT expansion relies on the assumption that the typical energy scale of the process remains below the cutoff scale $\Lambda$ [16, 8].
-In practice, the absence of events with extremely large $m_{\mu\mu bb}$ or jet transverse momenta indicates that the analysis is dominated by energy scales safely below the assumed new physics cutoff.
+Within the selected $Z$-pole region, the analysis is dominated by moderate kinematic scales set by the reconstructed invariant mass and object transverse momenta, which supports the use of the SMEFT description in the phase space considered.
+A simultaneous multi-operator fit is beyond the scope of the present exploratory study.
+Since correlations among operators can weaken individual constraints and may lead to approximate flat directions, the bounds quoted below represent individual one-operator-at-a-time expected constraints.
+A global treatment including correlations with other electroweak and flavor observables would be required for a complete SMEFT interpretation.
 III Event generation and simulation
 All processes are generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8 [17].
-Detector-level effects are simulated using Delphes [16] with an CMS-like detector configuration.
+Detector-level effects are simulated using Delphes [18] with a CMS-like detector configuration.
 Unless stated otherwise, proton–proton collisions at a center-of-mass energy of $\sqrt{s}=13~\mathrm{TeV}$ are assumed throughout this analysis.
-The signal process considered in this work is the four-body decay of the Z boson into a dimuon pair and a bottom–antibottom quark pair: $Z\to\mu\mu bb$.
-At leading order, the four-body final state arises predominantly from final-state radiation or off-shell electroweak boson exchange [7].
-Representative Feynman diagrams contributing to the signal process are shown in Fig.
-1.
+All numerical results presented below should be understood as leading-order constraints within the simulation setup described here.
+Higher-order QCD and electroweak corrections, as well as scale and PDF uncertainties, are not included in the baseline simulation.
+For several of the processes considered here, next-to-leading-order predictions can be obtained with public Monte Carlo tools and may modify both the normalization and the shapes of kinematic distributions.
+Their inclusion is left for future precision studies.
+The primary final state considered in this work is $pp\to\mu^{+}\mu^{-}b\bar{b}$ in the reconstructed $Z$-pole region.
+For brevity, we refer to this topology as the $Z\to\mu\mu b\bar{b}$ channel.
+At the amplitude level, however, the same final state may receive contributions from off-shell electroweak bosons and photon-mediated diagrams.
+The present analysis should therefore be understood as a fiducial sensitivity study of the $\mu^{+}\mu^{-}b\bar{b}$ final state around the $Z$ resonance rather than a fully gauge-invariant separation of individual diagram classes.
 Signal events are generated at parton level using MadGraph5_aMC@NLO.
 Basic generator-level cuts are applied to ensure numerical stability and reflect the kinematic acceptance of the detector: $p_{T}^{\ell}>10~\mathrm{GeV}$, $|\eta^{\ell}|<2.5$ for charged leptons, and $p_{T}^{q}>10~\mathrm{GeV}$, $|\eta^{q}|<2.5$ for quarks.
-The resulting leading-order cross section for the signal process is $\sigma(Z\to\mu^{+}\mu^{-}b\bar{b})=0.750~\mathrm{fb}$, where the quoted uncertainty corresponds to the Monte Carlo integration error.
+The resulting leading-order cross section for the selected signal topology is $\sigma=0.750~\mathrm{fb}$ at generator level before the full detector-level selection.
 The reconstructed kinematic distributions shown in Fig.
@@
 In Fig.
-2(a), the invariant mass distribution of the $b\bar{b}$ system is considerably broader than that of dimuon system, as expected from jet energy resolution effects and QCD radiation.
+2(a), the invariant mass distribution of the $b\bar{b}$ system is considerably broader than that of the dimuon system, as expected from jet energy resolution effects and QCD radiation.
 Nevertheless, the combined four-body invariant mass $m_{\mu\mu bb}$ displays a clear enhancement around $m_{Z}$, indicating that the signal topology can be reliably reconstructed at detector level despite the presence of hadronic final states.
@@
 ((b))
-Figure 2: Detector-level kinematic distributions for the process $Z\to\mu\mu bb$ obtained from Delphes simulation.
+Figure 2: Detector-level kinematic distributions for the selected $\mu^{+}\mu^{-}b\bar{b}$ final-state topology obtained from Delphes simulation.
 The dominant background arises from top-quark pair production $pp\to t\bar{t}$, with both top quarks decaying semileptonically, $t\to b\mu^{+}\nu_{\mu}$ and $\bar{t}\to\bar{b}\mu^{-}\bar{\nu}_{\mu}$ [15, 17].
@@
 Subleading backgrounds from electroweak processes involving photons are also considered, including $Z\gamma$, $\gamma\gamma$, and mixed $\gamma Z$ production, where the photon converts into a fermion pair.
+We note that electroweak amplitudes leading to the same $\mu^{+}\mu^{-}b\bar{b}$ final state can in principle interfere.
+In the present exploratory analysis, the samples are organized according to their dominant physical topology in order to estimate their relative impact after selection.
+A fully gauge-invariant treatment of the irreducible electroweak $\mu^{+}\mu^{-}b\bar{b}$ contribution, including all interferences among $Z^{\ast}$- and $\gamma^{\ast}$-mediated diagrams, is beyond the scope of this work.
+This point is taken into account by interpreting the numerical results as channel-specific constraints within the event-generation setup described above.
 A summary of all signal and background processes and their corresponding leading-order cross sections is given in TABLE 1.
 Process Cross section [pb]
-$pp\to Z,\ Z\to\mu^{+}\mu^{-}b\bar{b}$ $7.50\times 10^{-4}$
+$pp\to\mu^{+}\mu^{-}b\bar{b}$, $Z$-pole topology $7.50\times 10^{-4}$
 $pp\to t\bar{t}$, $t\to b\mu^{+}\nu_{\mu},\bar{t}\to\bar{b}\mu^{-}\bar{\nu}_{\mu}$ $4.38$
@@
 Table 1: Leading-order cross sections for the signal and background processes considered in this analysis.
-The quoted uncertainties correspond to Monte Carlo integration errors.
-The merged ZZ+jets sample, combining matrix-element calculations with up to two additional jets, is used in the final analysis, while the individual samples are shown for reference.
+The values are obtained at generator level before the full detector-level event selection and are intended to summarize the relative normalization of the simulated samples.
+The merged ZZ+jets sample, combining matrix-element calculations with up to two additional jets, is used in the final analysis.
 All generated parton-level events are passed to Pythia8 for parton showering and hadronization.
@@
 A jet is considered a $b$-jet if its $b$-tag discriminator exceeds a threshold of $0.5$, corresponding to a typical medium working point used in CMS analyses.
-Jets are identified as $b$-jets using the $b$-tagging discriminator provided by Delphes, adopting a medium working point.
 To emulate the trigger requirements of LHC experiments, events are required to satisfy at least one of the muon-based trigger conditions listed in TABLE 2.
@@
 To suppress backgrounds containing genuine missing transverse energy, such as top-quark pair production, events are required to have $\mathrm{MET}<30~\mathrm{GeV}$.
-This exploits the absence of neutrinos in the signal process $Z\to\mu\mu bb$.
+This exploits the absence of prompt neutrinos in the selected $\mu^{+}\mu^{-}b\bar{b}$ signal topology.
 Events are required to contain at least two reconstructed jets with $p_{T}>25~\mathrm{GeV}$, among which at least two must be identified as $b$-jets.
@@
 For events passing all selection requirements, the invariant mass of the $\mu\mu bb$ system is reconstructed using the four-momenta of the two leading isolated muons and the two selected $b$-jets.
-The reconstructed invariant mass $m_{\mu\mu bb}$, serves as the primary observable for signal extraction and the subsequent SMEFT analysis [7].
+The reconstructed invariant mass $m_{\mu\mu bb}$ serves as the primary observable for signal extraction and the subsequent SMEFT analysis [7].
 The signal region is defined as $80~\mathrm{GeV}<m_{\mu\mu bb}<115~\mathrm{GeV}$, centered around the $Z$ boson mass.
-This window retains the majority of the signal while efficiently suppressing non-resonant backgrounds.
-This window is chosen to maximize sensitivity to SMEFT-induced rate deviations while minimizing contamination from non-resonant backgrounds.
+This window retains a large fraction of the selected signal topology while reducing non-resonant backgrounds.
+It provides a simple and experimentally motivated signal region for the subsequent SMEFT constraint extraction.
 FIG.
@@
 FIG.
-3(b) focuses on the signal region defined by $80<m_{\mu\mu bb}<115$ GeV, which is centered around the $Z$ boson mass.
-In this window, the signal contribution becomes clearly distinguishable, while the $t\bar{t}$ background is strongly suppressed by the missing transverse energy requirement.
-The remaining backgrounds, including $ZA$ and $ZH$ production, exhibit similar kinematic features but are subleading in rate.
-This demonstrates that the chosen invariant mass window provides an optimal balance between signal efficiency and background rejection, forming a robust basis for the subsequent SMEFT analysis.
-The clear localization of the signal contribution within the $Z$-mass window demonstrates that the invariant mass $m_{\mu\mu bb}$ provides a powerful discriminator for isolating SMEFT effects in this channel.
-This observable is therefore chosen as the primary input for the subsequent reweighting and limit-setting procedure.
-V SMEFT Reweighting and Constraining
-The impact of dimension-six operators on the process $Z\to\mu\mu bb$ is evaluated within the SMEFT framework using an event reweighting technique [31, 32].
-In this approach, the effect of each WC on the cross section and kinematic distributions is obtained by reweighting SM Monte Carlo events.
-The total event weight includes the Standard Model contribution as well as linear and quadratic corrections from dimension-six operators, allowing an efficient exploration of the WCs parameter space without regenerating new event samples.
-SMEFT effects are incorporated using the SMEFTsim_general_MwScheme_UFO model from the SMEFTsim 3.0 framework [6, 22], which provides a complete and gauge-invariant implementation of dimension-six operators in the $M_{W}$ input scheme.
+3(b) focuses on the signal region defined by $80<m_{\mu\mu bb}<115$ GeV.
+In this window, the relative contribution of the selected signal topology is enhanced, while the $t\bar{t}$ background is reduced by the missing transverse energy requirement.
+The remaining electroweak and associated-production backgrounds, including $Z\gamma$ and $ZH$ production, exhibit similar kinematic features but are subleading in rate after the applied selection.
+The invariant mass $m_{\mu\mu bb}$ is therefore used as the main observable defining the signal region for the subsequent reweighting and statistics-only constraint extraction.
+V SMEFT Reweighting and Constraint Extraction
+The impact of dimension-six operators on the selected $\mu^{+}\mu^{-}b\bar{b}$ final-state topology is evaluated within the SMEFT framework using an event reweighting technique [26, 27].
+In this approach, the dependence of the selected event yield on each WC is obtained by reweighting Monte Carlo events.
+The total event weight includes the SM contribution as well as linear and quadratic corrections from dimension-six operators, allowing an efficient exploration of the WC parameter space without regenerating independent samples for every parameter point.
+SMEFT effects are incorporated using the SMEFTsim_general_MwScheme_UFO model from the SMEFTsim 3.0 framework [14], which provides a gauge-invariant implementation of the relevant dimension-six operators in the $M_{W}$ input scheme.
 The $M_{W}$ scheme is adopted to consistently account for electroweak input-parameter shifts induced by dimension-six operators.
-For each generated event, a set of precomputed weights corresponding to different values of $\theta$ is stored and propagated through parton showering, hadronization, and detector simulation, ensuring a consistent treatment of SM and SMEFT contributions [15, 16].
+For each generated event, a set of precomputed weights corresponding to different values of $\theta$ is stored at generator level.
+After parton showering, hadronization, detector simulation, and event selection, these weights are used to parametrize the SMEFT dependence of the selected yield.
 After applying the full event selection, the expected number of signal events is expressed as a quadratic function of the WCs parameter $\theta\equiv C_{i}/\Lambda^{2}$,
@@
 where $c$ denotes the Standard Model (SM) expectation, $b$ corresponds to the interference between the SM and dimension-six amplitudes, and $a$ encodes the pure quadratic SMEFT contribution [7].
-The coefficients $a$, $b$, and $c$ are determined by fitting the event yields obtained from reweighted samples generated at several fixed values of $\theta$.
+In principle, the reweighting information allows the SM contribution, the SM–dimension-six interference term, and the quadratic dimension-six contribution to be extracted separately.
+In the present analysis, the coefficients $a$, $b$, and $c$ are determined by fitting the reweighted event yields at several benchmark values of $\theta$.
+This provides a simple numerical implementation of the same quadratic dependence and is used as a consistency check of the reweighting behavior after event selection.
+The fitted parametrization is used only in the parameter region relevant for the quoted confidence intervals.
 Constraints on the WCs are derived using a likelihood-based statistical approach [28].
@@
 $q(\theta)=-2\ln\frac{\mathcal{L}(\theta)}{\mathcal{L}(0)}=2\left[N(\theta)\ln\frac{N(\theta)}{N_{\mathrm{SM}}}+N_{\mathrm{SM}}-N(\theta)\right],$ (5)
-which, under the assumptions of Wilks’ theorem, follows a $\chi^{2}$ distribution with one degree of freedom [33].
+which is compared with a $\chi^{2}$ distribution with one degree of freedom under the large-sample approximation of Wilks’ theorem [29].
 The 95% confidence level (C.L.) interval on $\theta$ is obtained by requiring $q(\theta)=3.84$.
-The resulting constraint is quoted both in terms of the WCs, $\theta_{95}=C_{i}/\Lambda^{2}$, and in terms of the relative deviation of the expected event yield with respect to the SM prediction,
-$R_{95}\equiv\frac{N(\theta_{95})}{N_{\mathrm{SM}}}.$ (5)
-This procedure is applied independently to each dimension-six operator, with all other WCs set to zero, providing a measure of the sensitivity of the $Z\to\mu\mu bb$ channel to physics beyond the Standard Model within the SMEFT framework.
+No nuisance parameters are included in the baseline likelihood.
+The resulting intervals therefore correspond to statistics-only expected constraints within the baseline likelihood setup.
+This procedure is applied independently to each dimension-six operator, with all other WCs set to zero, providing channel-specific constraints on possible SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state.
 In this analysis, a flavor-specific setup is adopted, in which the WCs are defined for individual fermion generations.
 In particular, operators involving second-generation leptons ($\ell_{2}$) and third-generation quarks ($q_{3}$) are considered explicitly.
-This choice allows a direct probe of flavor-resolved SMEFT effects in the $Z\to\mu\mu bb$ final state, beyond the flavor-universal assumptions commonly employed in global SMEFT fits.
+This choice provides channel-specific constraints on flavor-resolved SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state, beyond the flavor-universal assumptions commonly employed in global SMEFT fits.
 FIG.
@@
 For each operator, the ratio of the SMEFT prediction to the SM expectation is fitted with a quadratic function, reflecting the structure of the squared matrix element, which includes both interference and pure SMEFT contributions.
-The observed quadratic behavior confirms the validity of the reweighting procedure and illustrates the relative importance of linear and quadratic terms for different operator classes.
+The observed quadratic behavior is consistent with the expected structure of the squared matrix element when one Wilson coefficient is varied at a time.
+It also illustrates the relative importance of linear and quadratic terms for different operator classes.
 Operators modifying the ZZ–fermion couplings exhibit sizable interference with the SM amplitude, leading to asymmetric constraints around the SM point.
@@
 This behavior highlights the particular sensitivity of this channel to four-fermion operators involving bottom quarks.
-The resulting 95% C.L.
+The resulting expected 95% C.L.
 intervals on the WCs obtained from this procedure are summarized in Table 4.
-The constraints obtained in this analysis can be compared with existing global SMEFT interpretations from the CMS Collaboration [34].
-In the CMS combined analysis, stringent limits are reported on operators that modify the couplings of electroweak gauge bosons to fermions, such as $C_{H\ell}^{(1,3)}$ and $C_{Hq}^{(1,3)}$, under flavor-universal or top-specific assumptions.
-These limits, typically at the level of $\mathcal{O}(10^{-3})$$\mathcal{O}(10^{-2})~\mathrm{TeV}^{-2}$ at 95% C.L., are primarily driven by electroweak precision observables, Higgs measurements, and top-quark processes.
+These intervals are statistics-only, one-operator-at-a-time expected constraints obtained at leading order.
 ((a))
@@
 ((f))
-Figure 4: Quadratic SMEFT fits for the six operators under study.
-Each panel shows the fitted yield as a function of the WCs for the corresponding operator, with points representing simulated events and the red line indicating the quadratic fit.
-The legend shows the operator name following the SMEFTsim 3.0 conventions.
-In contrast, the present analysis focuses on flavor-resolved WCs involving second-generation leptons and third-generation quarks.
-The constraints derived for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ can therefore be directly interpreted as channel-specific probes of the same operator classes considered in the CMS global fit, but obtained from a single decay mode and a cut-based analysis.
-The constraints obtained here for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ are found to be of the same order of magnitude as those from the CMS global fit, despite being derived from a single decay channel and a cut-based analysis.
-For the four-fermion operators $C_{\ell q}^{(1,3),2233}$, which describe contact interactions between second-generation leptons and third-generation quarks, no direct constraints are currently available from global SMEFT fits, where such flavor-specific four-fermion operators are typically neglected or marginalized.
-The limits presented in this work therefore provide the first process-specific constraints on these operators from $Z$-boson decays.
-These results demonstrate the unique sensitivity of the $Z\to\mu\mu bb$ final state to four-fermion interactions involving bottom quarks, which are difficult to access in inclusive or flavor-universal analyses.
-The projected HL-LHC dataset, with an integrated luminosity of 3000 fb-1, further improves the constraints and providing a flavor-resolved probe complementary to existing CMS global analyses.
-In particular, the HL-LHC projections confirm the unique potential of the $Z\to\mu\mu bb$ channel to constrain four-fermion interactions involving bottom quarks.
+Figure 4: Dependence of the selected signal-region yield on individual Wilson coefficients in the parameter range relevant for the expected 95% C.L.
+intervals.
+The curves are obtained from the quadratic parametrization $N(\theta)=c+b\theta+a\theta^{2}$ using SMEFT reweighted samples.
+Only one coefficient is varied at a time, while all other coefficients are set to zero.
+The results include both interference and quadratic dimension-six contributions.
+A direct numerical comparison with existing global SMEFT fits should be made with care.
+Global analyses typically combine electroweak precision data, Higgs measurements, diboson production, top-quark observables, and differential LHC measurements.
+They may also adopt different flavor assumptions, operator subsets, input schemes, and SMEFT truncation prescriptions.
+In particular, bounds may differ significantly depending on whether they are obtained one operator at a time or after marginalization over a larger parameter space, and whether quadratic dimension-six terms are retained.
+For this reason, the results in Table 4 should be viewed as channel-specific expected constraints rather than direct replacements for global SMEFT constraints.
+Their main value is to illustrate that the $\mu^{+}\mu^{-}b\bar{b}$ final state has potential sensitivity to flavor-resolved operators involving second-generation leptons and third-generation quarks, especially four-fermion structures that are not always explicitly resolved in flavor-universal fits.
+The HL-LHC luminosity scenario illustrates the statistical improvement expected at larger datasets, while a full high-luminosity interpretation would also require a dedicated treatment of systematic uncertainties.
 $C$ 138 fb-1 3000 fb-1
@@
 $C_{H\ell}^{(3),22}$ [0.028, 0.025][-0.028,\;0.025] [0.006, 0.005][-0.006,\;0.005]
-Table 4: 95% C.L. allowed regions for $C/\Lambda^{2}$ [TeV-2].
+Table 4: 95% C.L.
+allowed regions for $C/\Lambda^{2}$ [TeV-2].
+The results are obtained at leading order using the fixed-width approximation and include both interference and quadratic dimension-six contributions.
+V.1 Uncertainty treatment
+The likelihood used in the baseline results includes only the Poisson statistical uncertainty of the expected event yield.
+Experimental systematic uncertainties, such as luminosity, muon reconstruction efficiency, $b$-tagging efficiency, jet energy scale, and background normalization uncertainties, are not included.
+The resulting intervals therefore represent statistics-only expected constraints in the baseline setup.
+A complete treatment of systematic uncertainties would require nuisance parameters constrained by auxiliary measurements. Schematically, for each bin or signal region one may write
+$N_{k}(\vec{C},\vec{\eta})=N_{k,\mathrm{sig}}(\vec{C})(1+\eta_{s})+\sum_{b}N_{k,b}(1+\eta_{b}),$ (6)
+with Gaussian constraints on the nuisance parameters,
+$\mathcal{L}(\vec{C},\vec{\eta})=\prod_{k}\mathrm{Pois}\left(n_{k}\mid N_{k}(\vec{C},\vec{\eta})\right)\prod_{j}\exp\left(-\frac{\eta_{j}^{2}}{2\sigma_{j}^{2}}\right).$ (7)
+Including such effects would generally weaken the projected bounds, especially for the HL-LHC luminosity scenario where statistical uncertainties become smaller.
+Since the present work is intended as a first constraint extraction, the incorporation of experimental systematics is left for future work.
+In addition, dimension-six effects are included only for the selected signal topology in the baseline reweighting procedure, while the background samples are kept at their SM predictions.
+Operators modifying ZZ–fermion couplings may also affect electroweak backgrounds such as $ZZ$ or ZZ+heavy-flavor production.
+A full SMEFT interpretation should include these effects consistently together with irreducible interferences among electroweak amplitudes.
 VI Conclusions
-In this work, we have presented a phenomenological study of the $Z\to\mu\mu bb$ final state as a probe of physics beyond the Standard Model within the framework of the Standard Model Effective Field Theory.
-The analysis is entirely based on Monte Carlo simulations and aims at assessing the sensitivity of this channel to a selected set of dimension-six SMEFT operators.
-Signal and relevant background processes were generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia, and a fast detector simulation with Delphes.
-A realistic event selection strategy was implemented, including trigger requirements, lepton isolation, $b$-jet identification, and kinematic separation cuts.
-After event reconstruction, the invariant mass distribution of the $\mu\mu bb$ system, together with transverse momentum spectra of the final-state objects, was used to validate the signal modeling and background suppression.
-To quantify possible deviations from the Standard Model, we employed the SMEFT reweighting technique using a UFO implementation of dimension-six operators.
-For each operator under consideration, event weights corresponding to multiple values of the WC were extracted at the generator level and propagated through the full analysis chain.
-The total event yield after selection was shown to follow a quadratic dependence on the WC, reflecting the interference and pure SMEFT contributions to the cross section.
-A likelihood-based statistical analysis was performed by constructing a test statistic based on the expected event yields.
-Assuming the Standard Model prediction as the null hypothesis, we derived 95% confidence level constraints on the WCs by identifying the parameter values corresponding to the critical value of the test statistic.
-This procedure provides a transparent and robust estimate of the sensitivity of the $\mu\mu bb$ channel to SMEFT effects, in the absence of experimental systematic uncertainties associated with real data.
-The results demonstrate that the $Z\to\mu\mu bb$ final state constitutes a viable and complementary probe of electroweak and Higgs-related effective operators.
-Although the present study is limited to leading-order simulations and an inclusive cut-based analysis, it establishes a solid baseline for future improvements.
-Possible extensions include the incorporation of higher-order corrections, a differential shape analysis, the inclusion of systematic uncertainties, and a global fit combining multiple final states.
-Overall, this work highlights the potential of precision studies of multi-body $Z$-boson decay final states within the SMEFT framework and demonstrates that mixed leptonic–hadronic channels can provide unique and complementary information on flavor-resolved new physics interactions.
+In this work, we have investigated the $\mu^{+}\mu^{-}b\bar{b}$ final state in the reconstructed $Z$-pole region as a probe of selected SMEFT operators.
+The analysis focuses on flavor-resolved interactions involving second-generation leptons and third-generation quarks, including both four-fermion operators and operators that modify effective ZZ–fermion couplings.
+Signal and relevant background processes were simulated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8, and fast detector simulation with Delphes.
+The event selection includes muon trigger requirements, lepton isolation, $b$-jet identification, missing-energy suppression, and angular separation cuts.
+The reconstructed invariant mass of the $\mu^{+}\mu^{-}b\bar{b}$ system is used to define a signal region around the $Z$ pole and to extract constraints on possible SMEFT effects.
+The dependence of the selected event yield on individual Wilson coefficients was obtained using the SMEFT reweighting procedure.
+For each operator, the yield was parametrized as a quadratic function of the Wilson coefficient, including both the SM–dimension-six interference and the quadratic dimension-six contribution.
+Using an Asimov likelihood-ratio method, we derived expected 95% C.L.
+constraints on the selected operators at integrated luminosities of $138~\mathrm{fb}^{-1}$ and $3000~\mathrm{fb}^{-1}$.
+The results show that the $\mu^{+}\mu^{-}b\bar{b}$ final state can provide useful channel-specific constraints on flavor-resolved SMEFT interactions, especially on operators involving muons and bottom quarks.
+These constraints are complementary to those obtained from inclusive electroweak measurements, purely leptonic $Z$ decays, and global SMEFT analyses.
+The projected HL-LHC luminosity further improves the reach, illustrating the potential value of including mixed leptonic–hadronic $Z$-pole final states in future SMEFT studies.
+The present analysis provides a baseline phenomenological study of this channel.
+Several refinements can be incorporated in future work, including higher-order QCD and electroweak corrections, scale and PDF uncertainties, experimental systematic uncertainties, correlations among multiple Wilson coefficients, SMEFT-induced modifications of the total $Z$ width, and a more complete treatment of electroweak interference effects in the irreducible $\mu^{+}\mu^{-}b\bar{b}$ final state.
+These extensions would allow the constraints from this channel to be embedded more directly into global SMEFT fits and precision electroweak interpretations.
 Acknowledgements.